OpenStudy (anonymous):

in 1980 median family income was about $16000 and in 2000 it was about $43000 waht will be the median family income in 2006

OpenStudy (anonymous):

1980=$16000 2000=$43000 2006= ? \[\frac{43000-16000}{20}=$1350 per year\] 6x1350= $8100 so 2006 = 43000+8100 = $51100